OpenStudy (anonymous):

If you have a 45.0 V battery connected to a 10.0 Ω resistor and a 5.00 Ω resistor in series, what will be the total power output in the circuit? A. 45.0 W B. 135 W C. 203 W D. 405 W

OpenStudy (michele_laino):

total resistance is: R= 10+ 5=...

OpenStudy (anonymous):

15!

OpenStudy (michele_laino):

ok! so current is: I=45/15=...

OpenStudy (anonymous):

3!

OpenStudy (michele_laino):

ok! Then the dissipated power is: W= I* V= 3*45=...watts

OpenStudy (anonymous):

135!so our solution is choice B? :O

OpenStudy (michele_laino):

yes! since it is the outputted power

OpenStudy (anonymous):

ooh okie! thank you!!:)

OpenStudy (michele_laino):

:)